# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@ 3.2.2 Gradient Manipulation Methods
While PCGrad [33] corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) [34] extends it to a more general and adaptive formulation. In each iteration, if $\phi_{ij}<\bar{\phi}_{ij}$, GradVac updates the corrected gradient $\hat{{\bm{g}}}_{i}$ as:
$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\left\|\hat{{\bm{g}}}_{i}\right\|\left(\bar{\phi}_{ij}\sqrt{1-\phi_{ij}^{2}}-\phi_{ij}\sqrt{1-\bar{\phi}_{ij}^{2}}\right)}{\left\|{\bm{g}}_{j}\right\|\sqrt{1-\bar{\phi}_{ij}^{2}}}{\bm{g}}_{j},$ (23)
where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$. $\bar{\phi}_{ij}$ is initialized to 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter. Note that GradVac simplifies to PCGrad when $\bar{\phi}_{ij}=0$.
@@
Gradient Sign Dropout (GradDrop) [35] considers that conflicts arise from differences in the sign of gradient values. Thus, a probabilistic masking procedure is proposed to preserve gradients with consistent signs during each update.
3.2.3 Practical Speedup Strategy
@@
Gradient balancing methods suffer from high computational and storage costs. Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t. the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$. This can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters. Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], Nash-MTL [27], and FairGrad [28]) need to solve an optimization problem to obtain the objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs. Hence, several strategies are proposed to alleviate this problem.
@@
Sener and Koltun [6] use feature-level gradients (i.e., gradients w.r.t. the representations from the last shared layer) to replace the parameter-level gradients (i.e., gradients w.r.t. the shared parameters ${\bm{\theta}}$). Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs. This strategy is also adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [30].
@@
Liu et al. [19] randomly select a subset of objectives to calculate the update direction in each iteration. Navon et al. [27] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iterations instead of updating in each iteration. Although this strategy speeds up training, they observe that it may cause a noticeable drop in performance. Liu et al. [32] consider optimizing the logarithm of the MGDA objective [6] and propose a speedup strategy. Specifically, when solving problem (13) via one-step gradient descent, ${\bm{\lambda}}$ is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$. Note that
$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$
$\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right]^{\top},$ (24)
@@
3.3 Discussions
Loss balancing and gradient balancing methods effectively mitigate conflicts among different objectives and achieve a Pareto-optimal solution, which is important for MOO scenarios such as multi-task learning. However, they are limited to finding a single Pareto-optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning. In Sections 4 and 5, we introduce methods to achieve more diverse and controllable solutions.
The computational cost of loss balancing methods (except MOML [13, 14], Auto-$\lambda$ [15], and FORUM [16]) is almost the same as that of EW since they only need one backpropagation process in each training iteration. However, gradient balancing methods are much more computationally expensive due to the need of $m$ backpropagation processes to obtain each objective gradient, where $m$ is the number of objectives. Although some speedup strategies are proposed to alleviate this problem, they may cause a drop in performance, as detailed in Section 3.2.3. Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, as detailed in Section 6.1.
Figure 4: Distributions of solutions obtained by using representative methods: PTML [36] and EPO [37, 38] are preference vector-based methods, while GradHV [47] and MOO-SVGD [48] are methods not using preference vectors. The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
4 Finding a Finite Set of Solutions
In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs. This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the Pareto set. We first discuss methods based on preference vectors (Section 4.1), followed by approaches that do not require the use of preference vectors (Section 4.2). Solutions obtained by some representative methods on LSMOP1 [84] are shown in Figure 4, with additional details and more results provided in Appendix B.
4.1 Methods Based on Preference Vectors
@@
Pareto Multi-Task Learning (PMTL) [36], inspired by the idea of decomposition-based MOO algorithms [85], incorporates preference vectors as constraints on the objectives. For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to the other preference vectors:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$
$~{}~{}\mathrm{s.t.}~{}~{}r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,\quad j\in[n].$ (25)
@@
Let ${\bm{R}}$ be the matrix containing gradients of the active constraints, i.e, ${\bm{R}}=\big{[}\nabla r^{(j)}({\bm{\theta}})\big{]}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ containing indices of the active constraints. Using a method akin to MGDA [6], PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$. The descent direction is:
${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$ (26)
@@
where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
${\bm{\lambda}},{\bm{\beta}}=$ $\operatorname*{arg\,min}_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},$
[…]
@@
A limitation of PMTL is its weak constraint that only requires the objective vector to lie within sectors, which lacks precise control over the position of Pareto solutions.
@@
Exact Pareto Optimal search (EPO) [37, 38] is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors. EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with ${\bm{\alpha}}$, achieving precise alignment and Pareto optimality. Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ be the $i$-th column of ${\bm{C}}$. Let ${\bm{a}}=[a_{1},\dots,a_{m}]^{\top}$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$. At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and ${\mathbb{J}}^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$. Based on these sets, it determines ${\bm{\lambda}}$ for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
${\bm{\lambda}}=$ $\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}~{}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
@@
$\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right.$ (30)
@@
This updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector. However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (30)).
@@
To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [39] considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem. Optimizing Tchebycheff function addresses the previously mentioned drawbacks of EPO. However, its convergence rate is relatively slow. Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the Preference-based MGDA (PMGDA) [40] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient.
@@
FERERO [41] captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint. It achieves fast convergence rates of $\mathcal{O}(\epsilon^{-1})$ for deterministic gradients and $\mathcal{O}(\epsilon^{-2})$ for stochastic gradients, where $\epsilon$ is the error tolerance.
@@
The methods discussed above rely on a fixed set of preference vectors. However, in real-world applications where the shape of the Pareto front is unknown, a predefined set of preference vectors may not always result in well-distributed solutions. To overcome this limitation, GMOOAR [42] formulates the problem as a bi-level optimization problem. In the upper level, preference vectors are optimized to maximize either the hypervolume or uniformity of the solutions. In the lower level, solutions are optimized based on these preference vectors. This approach enables the algorithm to dynamically adjust preference vectors based on the given optimization problem, ensuring the desired solution distribution. UMOD [43] introduces an approach that aims to maximize the minimum distance between solutions:
$\max\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (31)
@@
4.2 Methods without Using Preference Vectors
Another line of research, directly optimize for a diverse set of Pareto-optimal solutions, does not rely on a given set of preference vectors.

### v3

@@ 3.2.2. Gradient Manipulation Methods
While PCGrad (Yu et al., 2020a) corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) (Wang et al., 2021b) extends it to a more general and adaptive formulation. In each iteration, if $\phi_{ij}<\bar{\phi}_{ij}$, GradVac updates the corrected gradient $\hat{{\bm{g}}}_{i}$ as:
(23) $\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\left\|\hat{{\bm{g}}}_{i}\right\|\left(\bar{\phi}_{ij}\sqrt{1-\phi_{ij}^{2}}-\phi_{ij}\sqrt{1-\bar{\phi}_{ij}^{2}}\right)}{\left\|{\bm{g}}_{j}\right\|\sqrt{1-\bar{\phi}_{ij}^{2}}}{\bm{g}}_{j},$
@@
where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$. $\bar{\phi}_{ij}$ is initialized to 0 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter. Note that GradVac simplifies to PCGrad when $\bar{\phi}_{ij}=0$. Gradient Sign Dropout (GradDrop) (Chen et al., 2020) considers that conflicts arise from differences in the sign of gradient values. Thus, a probabilistic masking procedure is proposed to preserve gradients with consistent signs during each update.
3.2.3. Practical Speedup Strategy
@@
Gradient balancing methods suffer from high computational and storage costs. Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t. the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$. This can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters. Moreover, many gradient balancing methods (such as MGDA (Sener and Koltun, 2018), CAGrad (Liu et al., 2021b), Nash-MTL (Navon et al., 2022), and FairGrad (Ban and Ji, 2024)) need to solve an optimization problem to obtain the objective weight $\bm{\lambda}$} in each iteration, which also increases the computational and memory costs. Hence, several strategies are proposed to alleviate this problem.
@@
Sener and Koltun (Sener and Koltun, 2018) use feature-level gradients (i.e., gradients w.r.t. the representations from the last shared layer) to replace the parameter-level gradients (i.e., gradients w.r.t. the shared parameters $\bm{\theta}$}). Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs. This strategy is also adopted by some gradient balancing methods, such as IMTL-G (Liu et al., 2021a) and Aligned-MTL (Senushkin et al., 2023).
@@
Liu et al. (Liu et al., 2021b) randomly select a subset of objectives to calculate the update direction in each iteration. Navon et al. (Navon et al., 2022) propose to update the objective weight $\bm{\lambda}$} every $\tau$ iterations instead of updating in each iteration. Although this strategy speeds up training, they observe that it may cause a noticeable drop in performance. Liu et al. (Liu et al., 2023) consider optimizing the logarithm of the MGDA objective (Sener and Koltun, 2018) and propose a speedup strategy. Specifically, when solving problem (13) via one-step gradient descent, $\bm{\lambda}$} is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$. Note that
(24) $\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right$$^{\top}$,
@@
3.3. Discussions
Loss balancing and gradient balancing methods effectively mitigate conflicts among different objectives and achieve a Pareto-optimal solution, which is important for MOO scenarios such as multi-task learning. However, they are limited to finding a single Pareto-optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning. In Sections 4 and 5, we introduce methods to achieve more diverse and targeted solutions.
The computational cost of loss balancing methods (except MOML (Ye et al., 2021, 2024b), Auto-$\lambda$ (Liu et al., 2022), and FORUM (Ye et al., 2024a)) is almost the same as that of EW since they only need one backpropagation process in each training iteration. However, gradient balancing methods are much more computationally expensive due to the need of $m$ backpropagation processes to obtain each objective gradient, where $m$ is the number of objectives. While various speedup strategies can address this problem, they often come at the cost of performance, as detailed in Section 3.2.3. Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, detailed in Section 6.1.
Figure 5. Distributions of solutions obtained by using representative methods: PMTL (Lin et al., 2019b) and EPO (Mahapatra and Rajan, 2020, 2021) are preference vector-based methods, while GradHV (Wang et al., 2017) and MOO-SVGD (Liu et al., 2021d) are methods not using preference vectors. The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors. As can be seen, MGDA (Sener and Koltun, 2018) demonstrates poor diversity due to its lack of preference incorporation. While the preference vector-based method PTML (Lin et al., 2019b) achieves better alignment with the preference vector, it still falls short of exact alignment. In contrast, EPO (Mahapatra and Rajan, 2020, 2021) successfully achieves exact alignment. GradHV (Wang et al., 2017) and MOO-SVGD (Liu et al., 2021d) can also generate diverse solutions, despite they do not use a preference vector.
4. Finding a Finite Set of Solutions
In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs. This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the Pareto set. We first discuss methods based on preference vectors (Section 4.1), followed by approaches that do not require the use of preference vectors (Section 4.2). Figure 5 shows solutions obtained by some representative methods on LSMOP1 (Cheng et al., 2016) after 5000 iterations of each algorithm.
4.1. Methods Based on Preference Vectors
@@
Pareto Multi-Task Learning (PMTL) (Lin et al., 2019b), inspired by the idea of decomposition-based MOO algorithms (Zhang and Li, 2007), incorporates preference vectors as constraints on the objectives. For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to the other preference vectors:
(25) $\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},~~\mathrm{s.t.}~~r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,~~j\in[n].$
@@
Let $\bm{R}$} be the matrix containing gradients of the active constraints, i.e, ${\bm{R}}=\big{[}\nabla r^{(j)}({\bm{\theta}})\big{]}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ containing indices of the active constraints. Using a method akin to MGDA (Sener and Koltun, 2018), PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$. The descent direction is:
(26) ${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$
@@
where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the Jacobian matrix. $\bm{\lambda}$} and $\bm{\beta}$} are coefficients obtained by solving the following quadratic programming problem:
(27) $\min_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},~~\text{s.t.}~~{\bm{\lambda}}^{\top}\bm{1}+{\bm{\beta}}^{\top}\bm{1}=1,\;\lambda_{i}\geq 0,\;\beta_{j}\geq 0.$
@@
A limitation of PMTL is it can only constrain objective vectors within sectors, which lacks precise control over the position of Pareto solutions.
@@
Exact Pareto Optimal search (EPO) (Mahapatra and Rajan, 2020, 2021) is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors. EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with $\bm{\alpha}$}, achieving precise alignment and Pareto optimality. Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ be the $i$-th column of $\bm{C}$}. Let ${\bm{a}}=[a_{1},\dots,a_{m}]^{\top}$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$. At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and $\mathbb{J}$$^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$. Based on these sets, it determines $\bm{\lambda}$} for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
@@
(30) $\max_{{\bm{\lambda}}\in\Delta_{m-1}}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),~~\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right$.
@@
This updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector. However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem.
@@
To address these drawbacks, Weighted Chebyshev MGDA (WC-MGDA) (Momma et al., 2022) considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem. Optimizing Tchebycheff function addresses the previously mentioned drawbacks of EPO. However, its convergence rate is relatively slow. Two methods for improving the Tchebycheff scalarization include smooth Tchebycheff scalarization, which replaces the non-smooth max operation with a smooth approximation, and Preference-based MGDA (PMGDA), which only requires the update direction to have a negative inner product with the exact constraint gradient.
@@
FERERO (Chen et al., 2024a) captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint. It achieves fast convergence rates of $\mathcal{O}(\epsilon^{-1})$ for deterministic gradients and $\mathcal{O}(\epsilon^{-2})$ for stochastic gradients, where $\epsilon$ is the error tolerance.
@@
The methods discussed above rely on a fixed set of preference vectors. However, in real-world applications where the shape of the Pareto front is unknown, a predefined set of preference vectors may not always result in well-distributed solutions. To overcome this limitation, GMOOAR (Chen and Kwok, 2022) formulates the problem as a bi-level optimization problem. In the upper level, preference vectors are optimized to maximize either the hypervolume or uniformity of the solutions. In the lower level, solutions are optimized based on these preference vectors. This approach enables the algorithm to dynamically adjust preference vectors based on the given optimization problem, ensuring the desired solution distribution. UMOD (Zhang et al., 2024a) introduces an approach that aims to maximize the minimum distance between solutions:
(31) $\max_{{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS}}\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),$
@@
4.2. Methods without Using Preference Vectors